OpenStudy (anonymous):

10. Write mc005-1.jpg in scientific notation. A. mc005-2.jpg B. mc005-3.jpg C. mc005-4.jpg D. mc005-5.jpg

OpenStudy (anonymous):

o_o

OpenStudy (anonymous):

0.56*10^4 A 56*10^ B 0.56*10^3 C 5.6*10^3 56*10

OpenStudy (anonymous):

@HourglassMage

OpenStudy (anonymous):

sry it wont chow the numbers when I copy and paste

OpenStudy (anonymous):

0.56x10^4 is the question?

OpenStudy (anonymous):

show

OpenStudy (anonymous):

yeah, she wants to know how it should be written in scientific notation @aajugdar

OpenStudy (anonymous):

but it looks like it's already written in scientific notation - isn't it?

OpenStudy (anonymous):

well,its easy

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

yeah i am confused with options and the question but i will tell her how it is anyway

OpenStudy (anonymous):

well, if you have \[0.56*10^4\] then you can write it as \[0.56*10*10^3 = 5.6*10^3\]

OpenStudy (anonymous):

similarly you can write it as \[5.6*10*10^2 =56*10^2\]

OpenStudy (anonymous):

well right...but it's already in scientific notation. So what the hell are the looking for here?

OpenStudy (anonymous):

and you can also write is as \[56*10*10 = 560*10 = 5600\]

OpenStudy (anonymous):

I suppose it's in a simpler form in 56 x 10^2...

OpenStudy (anonymous):

so just look at options,see which of these forms is written and just answer it

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

I have no idea where she went - looks like she just abandoned this question.

OpenStudy (anonymous):

hey sry guys os was acting up

OpenStudy (anonymous):

it logged me out sry

OpenStudy (anonymous):

hi, so what were the options for the question?